OpenStudy (anonymous):

Howdy good people can you help me with this question, find the distance between points 6,16 and -1, 14

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

let me take for a spin real fast like

OpenStudy (anonymous):

would it be (7,9)?

hartnn (hartnn):

you're trying to find the 'distance' right ? the answer would be a number, not the co-ordinates. take it step by step, whats x1 -x2 = 6- (-1) = ... y1-y2 =16-14 = ...

OpenStudy (anonymous):

so would it be 16 units?

hartnn (hartnn):

how did u get the "16" ?

OpenStudy (anonymous):

oh uh seems that i have made a mistake would the numbers come out to 9?

hartnn (hartnn):

x1 -x2 = 6- (-1) = 7 y1-y2 =16-14 = 2 \(\Large \sqrt{7^2+2^2} =.... ? \)

OpenStudy (anonymous):

53!

hartnn (hartnn):

yes, \(\sqrt{53}\) :)

OpenStudy (anonymous):

thank you so much you are a kind soul sir or mam

hartnn (hartnn):

you're most welcome ^_^